# AugGen: Augmenting Task-Based Learning in Professional Creative Software with LLM-Generated Scaffolded UIs

### v1

@@ 3. AugGen
In this section, we present our method and implementation to generate task-specific scaffolded interfaces for professional creative software by surfacing relevant tools, embedding domain concepts, and mapping the scaffolded interface back to the native software. The method is grounded in principles of task-centric design, adaptive interfaces, and scaffolded learning, and implemented as a multi-stage pipeline based on LLM-assisted UI generation.
@@ 3.1. Method
Task-specific tool mapping. The method begins by extracting a set of tools required for the user’s task. Hidden or nested functions are surfaced to reduce learning barriers and interface complexity (C1). This is motivated by task-centric UI design principles (lafreniere2014task) to manage interface complexity and reduce user cognitive load (sweller2011cognitive).
Scaffolded interface construction. The method constructs an interface that organizes selected tools into workflow stages to offer task guidance and eliminate the need for users to switch between external tutorials and the software interface (C2). This is grounded in adaptive UI design principles (knowledge_navigator; jameson2007adaptive) to offer in-context, task-aware guidance through UI configuration.
Progressive disclosure. The scaffolded interface supports user-selectable levels of complexity, which enables beginners to start with fundamental functions and expand the interface as they gain proficiency. This aligns with the training-wheels interface model (carroll1984training) and intermodal guidelines (cockburn2014supporting) to manage the learning process and encourage exploration with user control.
Concept labeling. Each tool is labeled with domain terminology to connect actions, concepts, and workflow steps, supporting concept learning during task execution (C3). This is motivated by prior research on learning by doing (rieman1996field; novick2009micro) and vocabulary extension (cockburn2014supporting) to explicitly reinforce learning engagement while performing tasks.
@@ 3.2.1. Technical Pipeline
Here we present a technical pipeline to implement the scaffolded interface. In this work, we demonstrate an example that implements the scaffolded interface in Blender (version 3.6), a professional 3D creative software tool. The implementation pipeline involves using an LLM (GPT-4o) to assist with user task analysis, task-relevant tool selection, and UI code generation. This section outlines the technical pipeline and the complete implementation details for reproducibility can be found in Appendix A. Figure 4 shows this pipeline.
@@
Figure 4. Technical implementation pipeline built upon the AugGen method. (1) The pipeline takes a user’s task description in natural language (e.g., perform UV unwrapping) that informs LLM-assisted task analysis to reason workflow stages and relevant domain concepts (e.g., marking seams, unwrapping, checking and visualization). (2) These workflow stages subsequently guide LLM-supported tool selection and complexity assessment, resulting in a candidate tool list (e.g., tools to mark seams and unwrap UV map). (3) The tool list is fed into LLM-based code generation for the UI implementation. (4) The code is executed to generate a scaffolded interface in professional creative software (e.g., custom panel in Blender). Users (e.g., UI designers, software educators, end users) are allowed to step in to refine LLM-generated artifacts (e.g., workflow, tool list, code) to adjust the scaffolded interface.:
@@
Workflow analysis and concept identification. The technical pipeline starts with taking a user’s task description in natural language as the input. AugGen then assembles the workflow analysis and concept identification prompt (Appendix A.1) that guides the LLM to decompose the user task into sub-tasks, forming the task workflow, and identify relevant domain concepts, which are associated with each sub-task. In this process, the LLM needs to analyze the required steps to perform the user task and classify these steps into workflow stages, as well as extract relevant domain concepts and explain them in a straightforward, concise description.
@@
Tool selection and complexity assessment. With the workflow stages and relevant concepts, AugGen selects needed tools for each stage and maps them to the software functionality. For each workflow stage, AugGen uses the tool selection prompt with reference to the software manuals for the LLM to select tools, e.g., keyboard and mouse operations and UI interactions. As part of this process, AugGen assembles the complexity assessment prompt to instruct the LLM to suggest a complexity level for each tool (e.g., basic, intermediate, advanced) based on its functionality from software manuals (e.g., manuals in (blendermanuals)), and add these complexity levels to the tool description. Prompts for tool selection and complexity assessment are available in Appendix A.2.
@@
UI code generation and tool labeling. With the selected tools, AugGen generates the code to implement each tool’s functionality and renders the tools in a scaffolded interface. Specifically, AugGen first assembles the functionality code generation prompt for the LLM to generate the code that executes each tool’s functionality by referring to the software APIs (e.g., software APIs in (blenderapidocs)). Essentially, each keyboard and mouse operation or UI interaction is translated into an executable functionally-equivalent code block. Next, AugGen uses the UI code generation prompt to guide the LLM to generate the code that converts the functionality code execution into UI interaction, e.g., button click or drop-down selection, such that users can initiate the execution of each tool through the scaffolded interface. In the UI code generation process, AugGen provides the LLM with an example code that outlines the interface layout to structure tools based on their corresponding workflow stages and groups tools according to their complexity levels. Lastly, AugGen assembles the tool labeling prompt to ask the LLM to add concept description as tooltips for each tool. Prompts for code generation are available in Appendix A.3.
@@
Code execution and UI testing. In the last stage, AugGen executes the generated UI code in the professional creative software to display the scaffolded interface (e.g., a custom panel in Blender). With this interface, UI designers, software educators, and end users, can test and evaluate the UI or make optional adjustments. They can adjust the intermediate artifacts offered to them in text, including the workflow stages, domain concepts, and tool selections, as well as the generated UI code for UI customization. These adjustments lead to iterative executions of the technical pipeline for the user to obtain updated scaffolded interfaces. Each execution takes 1-2 minutes.
@@ 3.2.2. Generalizability
This technical pipeline can be used to accommodate different user tasks in the same and different software tools. Broader tasks are achievable by substituting a new task description, in natural language, as input to this pipeline. The pipeline will then analyze the task workflow, extract relevant domain concepts, select tools, and generate the UI code to create new scaffolded interfaces. Beyond a single software tool, this pipeline can be used to generate scaffolded interfaces in other software that provides a manual of its functionality for workflow analysis and tool selection, and programming APIs for UI code generation. Example software include Photoshop (photoshopapi), Maya (maya_api), AutoCAD (autocad_api), Unity (unity_api), and Grasshopper (grasshopperapi).
@@ 4. Illustrative Scenarios
To illustrate the user experience in using our scaffolded interfaces, we walk readers through how a user can complete a 3D modeling task and learn relevant domain concepts. We compare the experience with performing the same task using the default interface of a professional creative software tool (Blender).
@@
Experience using the default UI. Performing this task in the default UI (Figure 5) requires the user to navigate different menus and modes. The user enters “Edit Mode”, switches to “Edge Selection” mode, and then starts selecting edges on the 3D mesh. After selecting the edges, the user marks them as seams in a dedicated UV menu via “Ctrl+E”. Following it, they “right-click” within the editing region to open an unwrapping operator to unwrap the 3D mesh along the marked seams. The resulting UVs are viewed and refined in a separate UV Editor using tools from a toolbar.
@@
Figure 5. UV unwrapping using the default UI of Blender. Tools are located sparsely and sometimes hidden in keyboard shortcuts or mouse right-clicks. Domain concepts like “seams”, “unwrap”, and “UV islands” are opaque to users with limited explanation.:
Experience using the scaffolded UI. Using our scaffolded interface, as shown in Figure 8, to work on this task, the user starts at the Basic complexity level (Figure 6) and sees a structured task workflow. They first work through the Marking Seams stage. All the tools needed to select edges and mark seams are organized under this stage. Following the suggested flow, they move to Unwrapping & Editing and use the tools to unwrap the UV map and perform a simple visual check using tools in Checking & Visualization. Seeking to improve this initial result, they switch the interface’s complexity level to Intermediate (Figure 7) to utilize newly available seam selection aids and refine the UV layout. For final polishing and access to the complete toolset, they select the Advanced level (Figure 8) to reveal specialized tools. At this level, they can employ advanced operators for accurate adjustments and use analysis features to minimize distortion and verify the UV map quality.
Figure 6. UV unwrapping Basic level of the scaffolded UI.:
Figure 7. UV unwrapping Intermediate level of the scaffolded UI.:
@@
Throughout the process, the user is guided by the task workflow indicated by the interface without the need to switch back and forth between tutorials and their task. They can find task-relevant tools grouped under the corresponding workflow stage with less struggle remembering tool locations or the dedicated keyboard shortcuts and mouse clicks. When working on their task, they can control the tool complexity from basic to advanced based on their expertise and progress. They also get the chance to link their operations to relevant domain concepts and learn these concepts during task execution. Although the scaffolded UI is shown on a separate custom panel, they can connect each tool with the default software’s menu, keyboard shortcuts, and mouse operations through tool labels.
Figure 8. UV unwrapping using the scaffolded UI. Task-relevant tools are selected and grouped based on the workflow stages. Users can select complexity levels from basic, intermediate, to advanced for progressive tool disclosure. Hovering over each tool shows a tooltip explaining the related domain concepts. To connect the scaffolded interface with the native software environment, keyboard shortcuts, mouse clicks, and the location of each tool in the menus or toolbars are labeled.:
@@ 5. User Study Design
To evaluate our scaffolded interfaces, here we introduce the user study design with beginner (Study 1; Section 5.1) and expert users (Study 2; Section 5.2).
@@ 5.1. Study 1: Beginner Users
From the study with beginners, we explore the following research questions (RQs) to evaluate the effectiveness of the scaffolded interface compared to the baseline (default Blender interface) regarding task performance, concept learning, and software experience.
@@
  • RQ1: How does the scaffolded UI influence task performance? This question explores whether task-aware interfaces with relevant tools impact user-perceived task load, task performance, task completion time (efficiency), and user confidence.
@@
  • RQ2: How does the scaffolded UI affect the learning of domain concepts and software operations? This question evaluates how progressive tool disclosure, concept-integrated UI organization, and bridging to native software help users understand domain concepts, connect concepts to tools, and learn the native software operations.
@@
  • RQ3: How does the user experience differ with the scaffolded UI? This question investigates the overall user experience, including user-perceived ease of use, sense of support vs. constraint, entry barriers, and reliance on external help and documentation.
@@
Experiment design. We adopted a between-subjects design where each participant was randomly assigned to one of two interface conditions (default Blender interface or our scaffolded interface) and to one of two task conditions. Task 1 asked participants to perform UV unwrapping on a cube, a beginner-level texturing task, while Task 2 involved constructing a human walk cycle, a basic animation task. We chose these tasks because they represent two main task pipelines of 3D modeling and animation and require domain concept understanding. These tasks also accommodate adjustable complexity levels for us to evaluate with both beginner and expert users (Section 5.2). We provided all participants with the same cube for Task 1 and the same rigged human character ready for animation for Task 2 to ensure experiment consistency. The interfaces for Tasks 1 and 2 are shown in Figure 8 and Appendix B.1, respectively. To ensure consistent study instruction, we provided a study guide (Appendix B.2) that introduced the basics of UV unwrapping and animation, as well as publicly available tutorial videos, with durations of 4.5 minutes for Task 1 and 7 minutes for Task 2. Task 2 was more complex than Task 1 due to the longer completion time and more required steps.
@@
Participant recruitment. We recruited 32 English-speaking beginners (11 female, 19 male, 2 non-binary; mean age = 27.31, SD = 7.34) with no or limited 3D modeling and animation experience via Prolific (100% approval rate) (prolific). Participants were randomly assigned in a fully balanced design to one of two interface conditions, (1) Baseline (default Blender interface), or (2) Ours (scaffolded interface), to complete either Task 1 or Task 2, resulting in eight participants per task-interface combination.
@@
Measurements. We used four questionnaires with 7-point Likert scale responses. For RQ1, we used the NASA Task Load Index (TLX) (hart2006nasa) and a custom task performance questionnaire. For RQ2 and RQ3, we used custom questionnaires on concept learning and user interface experience. These custom questionnaires were designed based on broadly used standard questionnaires and an instructional model, including Software Usability Measurement Inventory (SUMI) (kirakowski1996software), Instructional Materials Motivation Survey (IMMS) (keller2010imms), and ARCS model of motivation (keller2009motivational). All the questionnaires are available in Appendix B.3.
@@
Procedure. We invited eligible participants to a Zoom meeting. Before the study started, we emailed a consent form (IRB protocol # anonymized), which outlined the study goal, procedure, duration, and data collection. After providing informed verbal consent, participants filled out a pre-study demographics questionnaire ($\sim$5 minutes). We then shared a study guide via a Google Slides link in Zoom chat, which included task instructions and tutorials. Once participants confirmed they had read the guide and accurately described the task, we shared our desktop screen, gave them remote control, and asked them to complete their assigned task. With consent, we recorded the desktop screen to record their task actions. Participants were encouraged to think aloud, and we took notes when observing them. Based on a pilot study with two beginners that helped us estimate task completion time, we gave the participants 20 minutes for Task 1 and 30 minutes for Task 2. After the task, participants were invited to complete the post-study questionnaires ($\sim$10 minutes). The study concluded with a payment of 10 USD to participants who worked on Task 1 for completing the $\sim$35 minute study and 15 USD to participants who worked on Task 2 for completing the $\sim$45 minute study.
@@
Data analysis. To evaluate the effects of interface and task on participant responses, we conducted Type II two-way analyses of variance (ANOVA) (seabold2010statsmodels) with Interface (Ours vs. Baseline) and Task (Task 1 vs. Task 2) as between-subjects factors to examine the main and interaction effects on our measurements. Following significant ANOVA effects, we performed post-hoc pairwise comparisons using Bonferroni-adjusted t-tests (Vallat2018pingouin) to identify specific differences.

### v2

@@ 3. TaskLens
In this section, we present our method and implementation to generate task-conditioned scaffolded interfaces in professional creative software to support software learners by surfacing relevant tools, embedding domain concepts, and mapping the scaffolded interface back to the native software. The method is grounded in principles of task-centric design, adaptive interfaces, and scaffolded learning, and is implemented as a multi-stage pipeline based on LLM-assisted UI generation.
@@ 3.1. Method
Task-specific tool mapping. The method begins by extracting a set of tools required for the user’s task. Hidden or nested functions are surfaced to reduce learning barriers and interface complexity (C1). This is motivated by task-centric UI design principles (Lafreniere et al., 2014) to manage interface complexity and reduce user cognitive load (Sweller, 2011).
Scaffolded interface construction. The method constructs an interface that organizes selected tools into workflow stages to offer task guidance and eliminate the need for users to switch between external tutorials and the software interface (C2). This is grounded in adaptive UI design principles (Apple, 1987; Jameson, 2007) to offer in-context, task-aware guidance through UI configuration.
Progressive disclosure. The scaffolded interface supports user-selectable levels of complexity, which enables beginners to start with fundamental functions and expand the interface as they gain proficiency. This aligns with the training-wheels interface model (Carroll and Carrithers, 1984) and intermodal guidelines (Cockburn et al., 2014) to manage the learning process and encourage exploration with user control.
Concept labeling. Each tool is labeled with domain terminology to connect actions, concepts, and workflow steps, supporting concept learning during task execution (C3). This is motivated by prior research on learning by doing (Rieman, 1996; Novick et al., 2009) and vocabulary extension (Cockburn et al., 2014) to explicitly reinforce learning engagement while performing tasks.
@@ 3.2.1. Technical Pipeline
Here we present a technical pipeline to implement the task-conditioned scaffolded interface. In this work, we demonstrate an example that implements the scaffolded interface in Blender (version 3.6), a professional 3D creative software tool. The implementation pipeline involves using an LLM to assist with user task analysis, task-relevant tool selection, and UI code generation. This section outlines the technical pipeline and the complete implementation details for reproducibility can be found in Appendix A. Figure 4 shows this pipeline.
@@
Figure 4. Technical implementation pipeline built upon the TaskLens method. (1) The pipeline takes a user’s task description in natural language (e.g., perform UV unwrapping) that informs LLM-assisted task analysis to reason workflow stages and relevant domain concepts (e.g., marking seams, unwrapping, checking and visualization). (2) These workflow stages subsequently guide LLM-supported tool selection and complexity assessment, resulting in a candidate tool list (e.g., tools to mark seams and unwrap UV map). (3) The tool list is fed into LLM-based code generation for the UI implementation. (4) The code is executed to generate a scaffolded interface in professional creative software (e.g., custom panel in Blender). Users (e.g., UI designers, software educators, end users) are allowed to step in to refine LLM-generated artifacts (e.g., workflow, tool list, code) to adjust the scaffolded interface.
This figure shows the technical pipeline. Starting from the left, the pipeline input is the user’s task, and it’s passed to the LLM to reason the task workflow. The workflow is then passed to the LLM to select relevant tools. The tool list is further fed to the LLM to generate code for UI implementation.
@@
Workflow analysis and concept identification. The technical pipeline starts with taking a user’s task description in natural language as the input. TaskLens then assembles the workflow analysis and concept identification prompt (Appendix A.1) that guides the LLM to decompose the user task into sub-tasks, forming the task workflow, and identify relevant domain concepts, which are associated with each sub-task. In this process, the LLM needs to analyze the required steps to perform the user task and classify these steps into workflow stages, as well as extract relevant domain concepts and explain them in a straightforward, concise description.
@@
Tool selection and complexity assessment. With the workflow stages and relevant concepts, TaskLens selects needed tools for each stage and maps them to the software functionality. For each workflow stage, TaskLens uses the tool selection prompt with reference to the software manuals for the LLM to select tools, e.g., keyboard and mouse operations and UI interactions. As part of this process, TaskLens assembles the complexity assessment prompt to instruct the LLM to suggest a complexity level for each tool (e.g., basic, intermediate, advanced) based on its functionality from software manuals (e.g., manuals in (Blender, 2025b)), and add these complexity levels to the tool description. Prompts for tool selection and complexity assessment are available in Appendix A.2.
@@
UI code generation and tool labeling. With the selected tools, TaskLens generates the code to implement each tool’s functionality and renders the tools in a scaffolded interface. Specifically, TaskLens first assembles the functionality code generation prompt for the LLM to generate the code that executes each tool’s functionality by referring to the software APIs (e.g., software APIs in (Blender, 2025a)). Essentially, each keyboard and mouse operation or UI interaction is translated into an executable functionally-equivalent code block. Next, TaskLens uses the UI code generation prompt to guide the LLM to generate the code that converts the functionality code execution into UI interaction, e.g., button click or drop-down selection, such that users can initiate the execution of each tool through the scaffolded interface. In the UI code generation process, TaskLens provides the LLM with an example code that outlines the script’s major components, including imported libraries, the interface layout (structures tools based on their corresponding workflow stages and groups tools according to their complexity levels), and main functions (link tools’ functionality implementation to UI interaction). Based on our experiments, the example code can help to reduce syntax errors in LLM-generated code, such as missing imports, random UI layout, or inactive UI elements. Lastly, TaskLens assembles the tool labeling prompt to ask the LLM to add concept description as tooltips for each tool. Prompts for code generation are available in Appendix A.3.
@@
Code execution and UI testing. In the last stage, TaskLens executes the generated UI code in the professional creative software to display the scaffolded interface (e.g., a custom panel in Blender). With this interface, UI designers, software educators, and end users, can test and evaluate the UI or make optional adjustments. They can adjust the intermediate artifacts offered to them in text, including the workflow stages, domain concepts, and tool selections, as well as the generated UI code for UI customization. Given the fact that LLM-generated outputs can be different from run to run even when the temperature parameter is set as zero (Yuan et al., 2025) and not always follow users’ task instructions (Zhou et al., 2023), the intermediate artifact inspection allows users to verify if the results are aligned with their requirements and make edits if needed. These user adjustments lead to iterative executions of the technical pipeline for the user to obtain updated scaffolded interfaces. Each execution takes 1-2 minutes.
3.2.2. Implementation Details
We implemented our pipeline using GPT-4o as the LLM. The LLM-generated UI code occasionally contained errors such as missing imports or unbound UI controls. Providing example code as references helped to reduce these occurrences, and when errors did arise, the pipeline propagated them back to the LLM for self-correction, typically resolving within one to two iterations. As LLM coding capabilities advance, we envision increasingly robust pipeline execution.
@@ 3.2.3. Generalizability
This technical pipeline can be used to accommodate different user tasks in the same and different software tools. Broader tasks are achievable by substituting a new task description, in natural language, as input to this pipeline. The pipeline will then analyze the task workflow, extract relevant domain concepts, select tools, and generate the UI code to create new scaffolded interfaces. Beyond a single software tool, this pipeline can be used to generate task-conditioned scaffolded interfaces in other software that provides a manual of its functionality for workflow analysis and tool selection, and programming APIs for UI code generation, such as UXP for Photoshop [^7] https://developer.adobe.com/photoshop/uxp/2022/, CUI for AutoCAD [^8] https://help.autodesk.com/view/OARX/2024/ENU/?guid=GUID-71554E76-8FD5-4853-82CD-3587764CBCAC, and OpenMayaUI for Maya [^9] https://help.autodesk.com/view/MAYAUL/2023/ENU/?guid=MAYA_API_REF_cpp_ref_group_open_maya_u_i_html. We further discuss the generalizability in Section 7.1 and demonstrate the implemented scaffolded UIs for variation tasks in Appendix D and for other software (Maya) in Appendix E.
@@ 4. Illustrative Scenarios
To illustrate the user experience in using our task-conditioned scaffolded interfaces, we walk readers through how a user can complete a 3D modeling task and learn relevant domain concepts. We compare the experience with performing the same task using the default interface of a professional creative software tool (Blender).
@@
Figure 5. Default Blender UI to perform UV unwrapping. Tools are located sparsely and sometimes hidden in keyboard shortcuts or right-clicks. Domain concepts, e.g., “seams”, “unwrap”, and “UV islands”, are opaque to beginners with limited explanation.
This figure shows the default Blender UI for UV unwrapping. Left is the UV editing region, which shows the UV map, and right is the 3D mesh operation region, which shows the 3D model.
Performing the task using the default UI. Performing this task in the default software UI of Blender (Figure 5) requires the user to navigate different menus and modes. The user enters “Edit Mode”, switches to “Edge Selection” mode, and then starts selecting edges on the 3D mesh. After selecting the edges, the user marks them as seams in a dedicated UV menu via “Ctrl+E”. Following it, they “right click” within the editing region to open an unwrapping operator to unwrap the 3D mesh along the marked seams. The resulting UVs are viewed and refined in a separate UV Editor using tools from a toolbar.
@@
Figure 6. Our task-conditioned scaffolded UI’s Basic level to perform UV unwrapping.
This figure shows the basic level interface. It displays the fundamental tools needed for UV unwrapping.
Performing the task using the task-conditioned scaffolded UI. Using our task-conditioned scaffolded interface, as shown in Figure 8, to work on this task, the user starts at the Basic complexity level (Figure 6) and sees a structured task workflow. They first work through the Marking Seams stage. All the tools needed to select edges and mark seams are organized under this stage. Following the suggested flow, they move to Unwrapping & Editing and use the relevant tools to unwrap the UV map and perform a simple visual check using tools in Checking & Visualization. Seeking to improve this initial result, they switch the interface’s complexity level to Intermediate (Figure 7) to utilize newly available seam selection aids and refine the UV layout. For final polishing and access to the complete toolset, they select the Advanced level (Figure 8) to reveal specialized tools. At this level, they can employ advanced operators for accurate adjustments and use analysis features to minimize distortion and verify the UV map quality.
Figure 7. Our task-conditioned scaffolded UI’s Intermediate level to perform UV unwrapping.
This figure shows the intermediate level interface. It displays the fundamental and some advanced tools that can be used for UV unwrapping.
@@
During the process, the user is guided by the task workflow indicated by the interface without needing to switch back and forth between tutorials and their task. They can find task-relevant tools grouped under the corresponding workflow stage with less struggle remembering tool locations or dedicated keyboard shortcuts and mouse clicks. When working on task, they can control the tool complexity from basic to advanced based on their expertise and progress. They also get the chance to link their operations to the relevant domain concepts and learn these concepts during task execution. Although the scaffolded UI is shown on a separate custom panel, they can connect each tool with the default software’s menu, keyboard shortcuts, and mouse operations through tool labels.
Figure 8. Our task-conditioned scaffolded UI generated to perform UV unwrapping. Task-relevant tools are selected and grouped based on the workflow stages. Users can select complexity levels from basic, intermediate, to advanced for progressive tool disclosure. Hovering over each tool shows a tooltip explaining the related domain concepts. To connect the scaffolded interface with the native software environment, keyboard shortcuts, mouse clicks, and the location of each tool in the menus or toolbars are labeled.
This figure shows the task-conditioned scaffolded UI for UV unwrapping. This scaffolded UI appears as an add-on panel on the right side in Blender’s UI.
@@ 5. User Study Design
To evaluate our task-conditioned scaffolded interfaces, here we introduce the user study design with beginners, who are the software learners (Study 1; Section 5.1), and with experts, who are advanced users developed their skills and knowledge through learning (Study 2; Section 5.2).
@@ 5.1. Study 1: Beginner Users
From the study with beginners (the software learners), we explore the following research questions (RQs) to evaluate the effectiveness of the task-conditioned scaffolded interface compared to the baseline (default Blender interface) regarding task performance, concept learning, and software experience.
@@
  • RQ1: How does the task-conditioned scaffolded UI influence task performance? This question explores whether task-aware interfaces with relevant tools impact user-perceived task load, task performance, task completion time (efficiency), and user confidence.
@@
  • RQ2: How does the task-conditioned scaffolded UI affect the learning of domain concepts and software operations? This question evaluates how progressive tool disclosure, concept-integrated UI organization, and bridging to native software help users understand domain concepts, connect concepts to tools, and learn the native software operations.
@@
  • RQ3: How does the user experience differ with the task-conditioned scaffolded UI? This question investigates the overall user experience, including user-perceived ease of use, sense of support vs. constraint, entry barriers, and reliance on external help and documentation.
@@
Experiment design. We adopted a between-subjects design where each participant was randomly assigned to one of two interface conditions (default Blender interface or our task-conditioned scaffolded interface) and to one of two task conditions. Task 1 asked participants to perform UV unwrapping on a cube, a beginner-level texturing task, while Task 2 involved constructing a human walk cycle, a basic animation task. We chose these tasks because they represent two main task pipelines of 3D modeling and animation, and require domain concept understanding. These tasks also accommodate adjustable complexity levels for us to evaluate with both beginner and expert users (Section 5.2). We provided all participants with the same cube for Task 1 and the same rigged human character ready for animation for Task 2 to ensure experiment consistency. The interfaces for Tasks 1 and 2 are shown in Figure 8 and Appendix B.1, respectively. To ensure consistent study instruction, we provided a study guide (Appendix B.2) that introduced the basics of UV unwrapping and animation, as well as publicly available tutorial videos, with durations of 4.5 minutes for Task 1 and 7 minutes for Task 2. Task 2 was more complex than Task 1 due to the longer completion time and more required steps.
@@
Participant recruitment. We recruited 32 English-speaking beginners (11 female, 19 male, 2 non-binary; mean age = 27.31, SD = 7.34) with no or limited 3D modeling and animation experience on Prolific [^10] https://www.prolific.com/ (100% approval rate). Participants were randomly assigned in a fully balanced design to one of two interface conditions, (1) Baseline (default Blender interface), or (2) Ours (task-conditioned scaffolded interface), to complete either Task 1 or Task 2, resulting in eight participants per task-interface combination.
@@
Measurements. We used four questionnaires with 7-point Likert scale responses. For RQ1, we used the NASA Task Load Index (TLX) (Hart, 2006) and a custom task performance questionnaire. For RQ2 and RQ3, we used custom questionnaires on concept learning and user interface experience. These custom questionnaires were designed based on broadly used standard questionnaires and an instructional model, including Software Usability Measurement Inventory (SUMI) (Kirakowski, 1996), Instructional Materials Motivation Survey (IMMS) (Keller, 2025), and ARCS model of motivation (Keller, 2009). All the questionnaires are available in Appendix B.3.
@@
Procedure. We invited eligible participants to a Zoom meeting. Before the study started, we emailed a consent form, which outlined the study goal, procedure, duration, and data collection. After providing informed verbal consent, participants filled out a pre-study demographics questionnaire ($\sim$5 minutes). We then shared a study guide via a Google Slides link in Zoom chat, which included task instructions and tutorials. Once participants confirmed they had read the guide and accurately described the task, we shared our desktop screen, gave them remote control, and asked them to complete their assigned task. With consent, we recorded the desktop screen to record their task actions. Participants were encouraged to think aloud, and we took notes when observing them. Based on a pilot study with two beginners that helped us estimate task completion time, we gave the participants 20 minutes for Task 1 and 30 minutes for Task 2. After the task, participants were invited to complete the post-study questionnaires ($\sim$10 minutes). The study concluded with a payment of 10 USD to participants who worked on Task 1 for completing the $\sim$35 minute study and 15 USD to participants who worked on Task 2 for completing the $\sim$45 minute study.
@@
Data analysis. To evaluate the effects of interface and task on participant responses, we conducted Type II two-way analyses of variance (ANOVA) (Seabold and Perktold, 2010) with Interface (Ours vs. Baseline) and Task (Task 1 vs. Task 2) as between-subjects factors to examine the main and interaction effects on our measurements. Following significant ANOVA effects, we performed post-hoc pairwise comparisons using Bonferroni-adjusted t-tests (Vallat, 2018) to identify specific differences.